# Learning Individual Behavior in Agent-Based Models with Graph Diffusion Networks

### v1

@@ Acknowledgments
The authors wish to thank Alberto Novati for his contribution to the early draft of the code for the original ABM of the predator-prey system. We also thank Daniele Grattarola for insightful early discussions that supported the initial development of this work.
Supplemental Material
@@ Appendix A Neural models and training details
In this section, we provide a detailed overview of the core components of the Graph Diffusion Network (GDN) and its methodology. We begin by introducing the diffusion process (A.1), which defines the diffusion process to be reversed in order to generate future agent states starting from a sample of Gaussian noise. Next, we detail the Graph Diffusion Network architecture (A.2) and its components. We then discuss the loss and optimization strategy (A.3), covering the training objectives and gradient flow between the diffusion model and graph components. Following this, we outline the generation algorithm (A.4), where the iterative denoising process generates future agent states. Finally, we provide a summary of the experimental resources and execution time (A.5), detailing computational requirements and execution time.
@@ A.2 Graph Diffusion Network architecture
In order to make the network more stable, all features are scaled. In particular, agent states $\mathbf{Z}_{t}^{(i)}$ can contain both numerical and categorical features. Numerical features are scaled to the interval $[-1,1]$. In our experiments, we scaled numerical features with a standard scaler. After generation, they are scaled back to their original domain and, for integer numerical features, a binning function is applied afterwards. Categorical features are one-hot encoded.
@@ A.3 Loss and optimization
$L(\phi,\omega)=\mathbb{E}_{i,t,\tau,\epsilon}\left[||\epsilon-\epsilon_{\phi}(\boldsymbol{\tilde{Z}}_{t+1}^{(i)}(\tau),\boldsymbol{c}_{t}^{(i)})||^{2}\right]$ (5)
@@ A.4 Generation
1:$\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}}\leftarrow Data$
2:$\mathbf{g_{t}}^{(i)}=f_{\omega}(\mathbf{Z}_{t}^{(i)},\bigoplus_{j\in N_{t}^{(i)}}(\mathbf{Z}_{t}^{(i)},\mathbf{Z}_{t}^{(j)}))$
3:$\tilde{\mathbf{Z}}_{t+1}^{(i)}(\tau_{max})\sim\mathcal{N}(\mathbf{0,I})$
@@
We set $\sigma_{\tau}=\sqrt{\frac{1-\bar{\alpha}_{\tau-1}}{1-\bar{\alpha}_{\tau}}\beta_{\tau}}$. This choice of $\sigma_{\tau}$ is optimal for deterministically set points [17], which is the case for some update rules in ABMs (e.g. happy agents in Schelling and dead agents in Predator-Prey). Alternatively, one can also choose $\sigma_{\tau}=\sqrt{\beta_{\tau}}$, which is more optimal for normally distributed points.
A.5 Experiments resources and execution time
@@
All experiments were run in a cloud based server with 15 vCores, 180 GB of RAM and an NVIDIA A100 80GB PCIe GPU. Execution times depend on the size of datasets and GPU occupancy by other processes. In our experiments, where the training datasets were made of $R=500$ ramifications over $T=10$ timesteps with 2048 agents in the system for Predator-Prey and 1950 agents for Schelling, the training time over 100 epochs typically lasted between 7 to 12 hours. Generating a simulation of 25 timesteps for the entire system of agents takes roughly between 10 to 12 seconds for both Predator-Prey and Schelling, less than 0.5 seconds per timestep.
@@ B.1 Schelling model
Algorithm 2 presents the pseudo-code of the ABM. Of particular interest are lines 15–25, which describe how agents relocate by searching for an empty cell on the grid. It is clear that this search process is not a simple draw from a probability distribution, as in the framework by Arya et al. [3], but a much more complex trial and error process. The pseudocode makes it clear that agents are more likely to relocate to nearby positions rather than distant ones, due to the way direction and distance are sampled. This spatial bias will be quantitatively confirmed in Figure 13.
@@
[…]
@@
6:        $\mathcal{N}_{t}^{(i)}\leftarrow\{\,j\in A:(x_{t}^{(j)},y_{t}^{(j)})\in\mathrm{Moore}(x_{t}^{(i)},y_{t}^{(i)})\}$triangleright Moore neighborhood
7:        $s\leftarrow|\{\,j\in\mathcal{N}_{t}^{(i)}:C^{(j)}=C^{(i)}\}|,\quad n\leftarrow|\mathcal{N}_{t}^{(i)}|$triangleright Same-color neighbors, all neighbors
8:        $r\leftarrow\begin{cases}s/n&n>0\\ 0&n=0\end{cases}$triangleright Similarity ratio
@@
19:           $\theta\sim\mathrm{Uniform}(0,2\pi),\;d\sim\mathrm{Uniform}(0,d_{\max})$triangleright Random direction, distance up to $d_{\max}$
20:           $\Delta x\leftarrow\lfloor d\cos\theta\rfloor,\;\Delta y\leftarrow\lfloor d\sin\theta\rfloor$triangleright Convert to grid movement
21:           $x^{*}\leftarrow(x_{t}^{(i)}+\Delta x)\bmod(L),\;y^{*}\leftarrow(y_{t}^{(i)}+\Delta y)\bmod(L)$triangleright Wrap around border
@@
23:               $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(x^{*},y^{*})$triangleright Move to new location
@@
29:        $Z_{t+1}^{(i)}\leftarrow Z_{t}^{(i)}$triangleright Happy agents stay put
@@ B.2 Predator-Prey model
The Predator-Prey ABM is a simulation model that captures the dynamics of interacting populations over time. We use a slightly adapted version of the model introduced in Ref. [39] (Algorithm 3 presents the pseudo-code of our ABM). Agents $i\in A$ occupy a two-dimensional grid of size $L\times L$. Each agent’s state at time $t$ is given by its kind (either Prey or Predator), its life phase (Unborn, Alive, Pregnant, or Dead), and its position on the grid:
@@
All transitions are governed by a matrix $\Psi$, which specifies deterministic rules as probabilities equal to 1, and defines stochastic transitions through probabilities that depend on spatial inter-species interactions. We report the values of the matrix $\Psi$ that define our four experimental settings in Tables 2-5. For instance, a Prey that interacts with a Predator is more likely to die compared to a Prey that does not interact with a Predator. Similarly, a Predator that does not interact with a Prey is more likely to die compared to a Predator that does interact with a Prey, as it is more likely to starve. These spatial interactions are defined by proximity: an agent interacts with others located in its Von Neumann neighborhood (i.e., the 4-neighborhood).
@@
[…]
6:        $\text{parent}(i)\sim\text{Uniform}(\{j\in A:j\neq i\wedge K^{(j)}=K^{(i)}\})$triangleright Assign parent
@@
8:    $\mathbf{Z}_{0}^{{}^{(i)}}=(K^{(i)},f^{(i)}_{0},x_{0}^{(i)},y^{(i)}_{0})$triangleright Agent state
@@
14:               $\mathcal{N}_{t}^{(i)}\leftarrow\mathrm{VonNeumann}(x_{t}^{(i)},y_{t}^{(i)})$triangleright Von Neumann neighborhood
[…]
18:                  $(x_{t+1}^{(i)},y_{t+1}^{(i)})\sim\mathrm{Uniform}(\mathcal{N}_{t}^{(i)})$triangleright Move randomly
[…]
20:                  $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(\emptyset,\emptyset)$triangleright Agent dies
[…]
22:                  $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(x_{t}^{(i)},y_{t}^{(i)})$triangleright Stay in place
@@
[…]
26:               $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(x_{t}^{(i)},y_{t}^{(i)})$triangleright Stay in place
@@
[…]
@@
31:               $j\leftarrow\mathrm{parent}(i)$triangleright Get parent
[…]
@@
35:                  $(x_{t+1}^{(i)},y_{t+1}^{(i)})\sim\mathrm{Uniform}(\mathcal{N}_{t}^{(j)})$triangleright Place near parent
@@
[…]
@@
[…]
@@ Appendix C Further experimental details and results
In this section, we provide additional details on the experiments and present further results. In the subsection on experimental design (C.1), we describe the micro and macro metrics used, including the number of points over which these metrics are computed. We then present additional qualitative results on reproducing emergent segregation in the Schelling model (C.2) and emergent oscillations in predator-prey dynamics (C.3). Finally, we provide further quantitative results for both models (C.4).

### v2

@@ Acknowledgments
The authors wish to thank Daniele Grattarola and Federico Cinus for insightful early discussions that supported the initial development of this work. We also thank Alberto Novati for his contribution to the early draft of the code for the original ABM of the predator-prey system.
Learning Individual Behavior in Agent-Based Models with Graph Diffusion Networks Supplementary Material
@@ Appendix A Neural models and training details
In this section, we provide a detailed overview of the core components of the Graph Diffusion Network (GDN) and its methodology. We begin by introducing the diffusion process (A.1), which defines the diffusion process to be reversed to generate future agent states starting from a sample of Gaussian noise. Next, we detail the Graph Diffusion Network architecture (A.2) and its components. We then discuss the loss and optimization strategy (A.3), covering the training objectives and gradient flow between the diffusion model and graph components. Following this, we outline the generation algorithm (A.4), where the iterative denoising process generates future agent states. Finally, we provide a discussion on the computational costs and scalability of our methodology (A.5).
@@ A.2 Graph Diffusion Network architecture
To make the network more stable, all features are scaled. In particular, agent states $\mathbf{Z}_{t}^{(i)}$ can contain both numerical and categorical features. Numerical features are scaled to the interval [1,1][-1,1]. In our experiments, we scaled numerical features with a standard scaler. After generation, they are scaled back to their original domain and, for integer numerical features, a binning function is applied afterwards. Categorical features are one-hot encoded.
@@
GNN-only ablated model
The only architectural change of the GNN-only ablated model is in the MLP following the aggregation pass of the GNN: its output dimension is set to the agent-state dimension rather than the embedding dimension. While the original GNN uses hidden layers of size [32, 64, 128], the ablated model adopts a symmetrical structure: [32, 64, 128, 128, 64, 32]. We trained for 100 epochs with the Adam optimizer (learning rate $2\times 10^{-5}$) and a mini-batch size of 16.
@@ A.3 Loss and optimization
[…]
@@ A.4 Generation
[…]
@@
We set $\sigma_{\tau}=\sqrt{\frac{1-\bar{\alpha}_{\tau-1}}{1-\bar{\alpha}_{\tau}}\beta_{\tau}}$. This choice of $\sigma_{\tau}$ is optimal for deterministically set points [ddpm_original_paper], which is the case for some update rules in ABMs (e.g. happy agents in Schelling and dead agents in Predator-Prey). Alternatively, one can also choose $\sigma_{\tau}=\sqrt{\beta_{\tau}}$, which is more optimal for normally distributed points.
A.5 Computational costs and scalability.
In the following subsection, we provide additional details and discussion on the computational costs and scalability of our proposed methodology. In particular, details on the hardware employed to run our experiments, performance details with respect to the number of ramifications provided during training, as well as a discussion over the scalability of Graph Diffusion Networks to larger systems of agents, and on the interplay between the number of agents and the number of ramifications.
Runtime characteristics.
@@
All experiments were run in a cloud-based server with 15 vCores, 180 GB of RAM, and an NVIDIA A100 80GB PCIe GPU. Execution times depend on the size of the datasets and GPU occupancy by other processes. In our experiments, where the training datasets were made of $R=500$ ramifications over $T=10$ timesteps with 2048 agents in the system for Predator-Prey and 1950 agents for Schelling, the training time over 100 epochs typically lasted around 1 hour (around 37 seconds per epoch). Generating a simulation of 25 timesteps for the entire system of agents takes roughly 7.5 seconds for both Predator-Prey and Schelling, around 0.3 seconds per timestep.
Scalability with respect to number of ramifications.
During training, each training step processes the entire system of agents for one timestep and one ramification. Therefore, the runtime grows linearly with respect to the number of ramifications or timesteps. We present results from 10 experiment runs trained on one of the Predator-Prey datasets of parameter $\Psi_{1}$, to compare training time, micro-level metric (EMD), and macro-level metric (sMAPE) as the number of ramifications provided during training increases in Figure 5 and Table 2. As expected, training time increases linearly with the number of ramifications included in the dataset used to train the model. Furthermore, micro-level metrics decrease as the number of ramifications increases, showcasing the benefit of providing a higher number of stochastic outcomes to the model during training to better fit the stochastic rules that govern the original ABM. Macro-level metrics are subject to higher variance and do not show a trend as clear as with the micro-level metrics, still the best results are reached with the highest number of ramifications.
Scalability with respect to number of agents.
The experiments covered in Section 4.2 are performed on mid-sized ABMs, with the total agent count ranging in the thousands (for Schelling there are 1950 agents in total and 2048 for Predator-Prey). Increasing the number of agents in the system (for example by increasing the grid size and keeping the density of agents constant) increases the training time sub-linearly. In fact, with mid-sized systems, we can train the network with batches that correspond to the whole set of agents. Thus, the number of training iterations per epoch does not change, but rather the size of the batch in input to the network. It should be noted that increasing the number of agents also increases the inference time (time required by the trained surrogate to produce a simulation for the whole system). We present results from 5 experiment runs trained on one of the Predator-Prey datasets of parameter $\Psi_{1}$, to compare training time, inference time, micro-level metric (EMD) and macro-level metric (sMAPE) as the total number of agents in the system increases by increasing grid size and keeping agent density constant in Figure 6 and Table 3. From our results, it can be noted that micro-level metrics slightly decrease as agent count increases, as well as macro-level metrics reach lower values in some of the experiments with higher agent count. We argue that increasing the number of agents in the system naturally increases the number of agent transitions available in the dataset, expanding the statistical coverage provided to the model to learn, and yielding slightly better results.
Scalability interplay between number of agents and ramifications.
Generating hundreds of futures per state can be costly for many-agent or long-horizon systems, but larger systems naturally provide more independent samples of similar conditions; thus, fewer ramifications can suffice for comparable statistical coverage. To evaluate the trade-off between data generation cost and predictive quality, we performed 5 experiment runs and trained our model on one Predator–Prey dataset (parameter set $\Psi_{1}$) with increasing agent counts and decreasing ramifications, keeping the training time per epoch approximately constant, and the total number of agent transitions roughly similar. All other hyperparameters match the main experiments. We observe from Table 4 that the errors (both sMAPE and EMD) remain close to the original results even when the number of agents grows substantially, keeping training time low thanks to a lower number of ramifications needed.
Table 2: Effect of ramifications $R$ on training time and performance.
Ramifications Time/epoch sMAPE EMD
50 $3.71\pm 0.02$ s $076\pm 0.019 0.0$$0161\pm 0.0003$
100 $7.44\pm 0.02$ s $071\pm 0.015 0.0$$0156\pm 0.0004$
250 $18.61\pm 0.06$ s $083\pm 0.039 0.0$$0115\pm 0.011$
500 $37.17\pm 0.12$ s $062\pm 0.030 0.0$$0088\pm 0.0007$
Table 3: Scaling with grid size, number of agents with fixed number of ramifications ($R=500$).
*: Mean time to produce a simulation of the whole system for $t=25$ timesteps.
Grid size Agents Time/epoch Inference time* sMAPE EMD
64 8192 $108.58\pm 0.30$ s $35.32\pm 2.03$ s $051\pm 0.026 0.0$$0082\pm 0.0003$
48 4608 $63.78\pm 0.32$ s $16.45\pm 0.72$ s $053\pm 0.026 0.0$$0083\pm 0.0005$
32 2048 $37.20\pm 0.12$ s $7.44\pm 0.01$ s $065\pm 0.033 0.0$$0092\pm 0.0007$
Table 4: Scaling with grid size, number of agents, and ramifications while keeping training time and transitions roughly constant.
Grid size Agents Ramifications Time/epoch sMAPE EMD
64 8192 180 $38.88\pm 0.18$ s $066\pm 0.027 0.0$$0095\pm 0.0005$
48 4608 300 $38.41\pm 0.25$ s $073\pm 0.024 0.0$$0092\pm 0.0005$
32 2048 500 $37.20\pm 0.12$ s $065\pm 0.033 0.0$$0092\pm 0.0007$
Figure 5: Training time, micro and macro metrics with respect to the number of ramifications provided during training for 10 experiments on one dataset of Predator-Prey with parameter set $\Psi_{1}$. Points indicate the mean value and error bars standard deviation over the 10 experiments.
Figure 6: Training time, micro and macro metrics with respect to the total number of agents in the system for 5 experiments on one dataset of Predator-Prey with parameter set $\Psi_{1}$. Points indicate the mean value and error bars standard deviation over the 5 experiments.
Figure 7: Training time, micro and macro metrics with increasing the total number of agents in the system and decreasing the number of ramifications to yield similar training times. Results are shown for 5 experiments on one dataset of Predator-Prey with parameter set $\Psi_{1}$. Points indicate the mean value and error bars standard deviation over the 5 experiments.
@@ B.1 Schelling model
Algorithm 2 presents the pseudo-code of the ABM. Of particular interest are lines 15–25, which describe how agents relocate by searching for an empty cell on the grid. It is clear that this search process is not a simple draw from a probability distribution, as in the framework by Arya et al. [arya2022automatic], but a much more complex trial and error process. The pseudocode makes it clear that agents are more likely to relocate to nearby positions rather than distant ones, due to the way direction and distance are sampled. This spatial bias will be quantitatively confirmed in Figure 16.
@@
4:  unhappy $\leftarrow\emptyset$ $\triangleright$ Reset unhappy list
@@
6:   $\mathcal{N}_{t}^{(i)}\leftarrow\{\,j\in A:(x_{t}^{(j)},y_{t}^{(j)})\in\mathrm{Moore}(x_{t}^{(i)},y_{t}^{(i)})\}$ $\triangleright$ Moore neighborhood
7:   $s\leftarrow|\{\,j\in\mathcal{N}_{t}^{(i)}:C^{(j)}=C^{(i)}\}|,\quad n\leftarrow|\mathcal{N}_{t}^{(i)}|$ $\triangleright$ Same-color neighbors, all neighbors
8:   $r\leftarrow\begin{cases}s/n&n>0\\ 0&n=0\end{cases}$ $\triangleright$ Similarity ratio
@@
19:     $\theta\sim\mathrm{Uniform}(0,2\pi),\;d\sim\mathrm{Uniform}(0,d_{\max})$ $\triangleright$ Random direction, distance up to $d_{\max}$
20:     $\Delta x\leftarrow\lfloor d\cos\theta\rfloor,\;\Delta y\leftarrow\lfloor d\sin\theta\rfloor$ $\triangleright$ Convert to grid movement
21:     $x^{*}\leftarrow(x_{t}^{(i)}+\Delta x)\bmod(L),\;y^{*}\leftarrow(y_{t}^{(i)}+\Delta y)\bmod(L)$ $\triangleright$ Wrap around border
@@
23:      $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(x^{*},y^{*})$ $\triangleright$ Move to new location
@@
29:   $Z_{t+1}^{(i)}\leftarrow Z_{t}^{(i)}$ $\triangleright$ Happy agents stay put
@@ B.2 Predator-Prey model
The Predator-Prey ABM is a simulation model that captures the dynamics of interacting populations over time. We use a slightly adapted version of the model introduced in Ref. [tang2022data] (Algorithm 3 presents the pseudo-code of our ABM). Agents $i\in A$ occupy a two-dimensional grid of size $L\times L$. Each agent’s state at time $t$ is given by its kind (either Prey or Predator), its life phase (Unborn, Alive, Pregnant, or Dead), and its position on the grid:
@@
All transitions are governed by a matrix $\Psi$, which specifies deterministic rules as probabilities equal to 1, and defines stochastic transitions through probabilities that depend on spatial inter-species interactions. We report the values of the matrix $\Psi$ that define our four experimental settings in Tables 5-8. For instance, a Prey that interacts with a Predator is more likely to die compared to a Prey that does not interact with a Predator. Similarly, a Predator that does not interact with a Prey is more likely to die compared to a Predator that does interact with a Prey, as it is more likely to starve. These spatial interactions are defined by proximity: an agent interacts with others located in its Von Neumann neighborhood (i.e., the 4-neighborhood).
@@
3:  $K^{(i)}\sim\text{Uniform}(\{\mathit{Prey},\mathit{Predator}\})$ $\triangleright$ Assign kind
4:  $f^{(i)}_{0}\sim\text{Uniform}(\{\mathit{Alive},\mathit{Unborn}\})$ $\triangleright$ Initial phase
5:  if $f_{0}^{(i)}=\mathit{Unborn}$ then
6:   $\text{parent}(i)\sim\text{Uniform}(\{j\in A:j\neq i\wedge K^{(j)}=K^{(i)}\})$ $\triangleright$ Assign parent
@@
8:  $\mathbf{Z}_{0}^{{}^{(i)}}=(K^{(i)},f^{(i)}_{0},x_{0}^{(i)},y^{(i)}_{0}) \triangleright$ Agent state
@@
14:      $\mathcal{N}_{t}^{(i)}\leftarrow\mathrm{VonNeumann}(x_{t}^{(i)},y_{t}^{(i)})$ $\triangleright$ Von Neumann neighborhood
15:      $\mathit{neighbor}\leftarrow\exists j\in A:\,(x_{t}^{(j)},y_{t}^{(j)})\in\mathcal{N}_{t}^{(i)}\wedge K^{(j)}\neq K^{(i)}$ $\triangleright$ Find opposite-kind neighbors
16:      $f_{t+1}^{(i)}\sim\mathrm{Categorical}\bigl(\Psi(K^{(i)},f_{t}^{(i)},\mathit{neighbor})\bigr)$ $\triangleright$ Random life phase update
17:      if $f_{t+1}^{(i)}=\mathit{Alive}$ then
18:       $(x_{t+1}^{(i)},y_{t+1}^{(i)})\sim\mathrm{Uniform}(\mathcal{N}_{t}^{(i)})$ $\triangleright$ Move randomly
19:      else if $f_{t+1}^{(i)}=\mathit{Dead}$ then
20:       $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(\emptyset,\emptyset)$ $\triangleright$ Agent dies
21:      else if $f_{t+1}^{(i)}=\mathit{Pregnant}$ then
22:       $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(x_{t}^{(i)},y_{t}^{(i)})$ $\triangleright$ Stay in place
@@
25:      $f_{t+1}^{(i)}\leftarrow\mathit{Alive}$
26:      $(x_{t+1}^{(i)},y_{t+1}^{(i)})\leftarrow(x_{t}^{(i)},y_{t}^{(i)})$ $\triangleright$ Stay in place
@@
28:      $f_{t+1}^{(i)}\leftarrow\mathit{Dead}$ $\triangleright$ Remain dead
@@
31:      $j\leftarrow\mathrm{parent}(i)$ $\triangleright$ Get parent
32:      if $f_{t}^{(j)}=\mathit{Pregnant}$ then
33:       $f_{t+1}^{(i)}\leftarrow\mathit{Alive}$ $\triangleright$ Born if parent is pregnant
@@
35:       $(x_{t+1}^{(i)},y_{t+1}^{(i)})\sim\mathrm{Uniform}(\mathcal{N}_{t}^{(j)})$ $\triangleright$ Place near parent
@@
37:       $f_{t+1}^{(i)}\leftarrow\mathit{Unborn}$ $\triangleright$ Remain unborn
@@
41:  if $\forall a\in A,\,f_{t}^{(a)}\in\{\mathit{Dead},\mathit{Unborn}\}$ then
@@ Appendix C Further experimental details and results
In this section, we provide additional details on the experiments and present further results. In the subsection on experimental design (C.1), we describe the micro and macro metrics used, including the number of points over which these metrics are computed. We then present additional qualitative results on reproducing emergent segregation in the Schelling model (C.2) and emergent oscillations in predator-prey dynamics (C.3). Then, we provide further quantitative results for both models (C.4). Finally, we provide a discussion on a macro-level baseline (C.5), specifically an autoregressive model of order one, or AR(1), a standard time-series model.